Two tangents are drawn to the circle C, such that they meet each other at a point outside the circle making an angle of x degrees. If the angle subtended by the two points of contact at the centre of the centre of the circle is 65 degrees, what is x?
Answer by josgarithmetic(39617)

This forms a quadrilateral with angle measures 65, t, t, and x.
x is outside of the circle, 65 is at the circle center, and two angles of t each are opposite from each other.

TANGENT, means the angle made ON the circle is 90 degrees. t=90 degrees.

65+90+90+x=360, because quadrilateral has 360 degrees for sum of interior angle measures.
x=360-65-180
x=180-65
x=115